OpenStudy (reverdeparadis):

Solve using substitution: y - 1/2x = 6 2x + 6y = 19

OpenStudy (anonymous):

For the first equation, add 1/2x to both sides: \[y=6+1/2x\] Substitute in the second equation for y: \[2x+6(6+1/2x)=19\] Distribute: \[2x+36+3x=19\] Add like terms: \[5x+36=19\] Subtract 36 from both sides: \[5x=-17\] Divide both sides by 5: \[x=-3.4\]